1) Supply and demand

P = 0.5QS + 30
P = -0.4QD + 120

a) Given the above equations, produce a chart illustrating both the supply and demand schedules in increments of 5 ranging from price = 50 to price = 110.
b) Solve for the equilibrium price and quantity and show your work.
c) Graph the result, labeling the axes, the supply and demand curves, the equilibrium point, and the price and quantity amounts. Use a proper scale.

d) Redo parts (a) through (c), this time reflecting a change in the supply curve to P = 0.5QS + 7.5.
e) Give two possible explanations for what could have caused such a change.
f) Now take the new graph (from part d) and assume that the government institutes a price floor equal to P = 80. Describe in detail what would be the impact on this market.
Answer #1

1, Given, P = 0.5Qs + 30

or, Qs = 2P - 60

P = -0.4Qd +120

Qp = 300 - 2.5P

a. Demand and Supply Schedule

Price QD QS $ 50 175 $ 55 162.5 $ 60 150 65 137.5 70 125 $ 75 112.5 80 100 100 87.5 110 90 75 120 95 62.5 130 $ 100 50 140 $

b. Equilibrium will be at that piont where Quantity demanded is equal to quantity supplied.

300 -2.5P = 2P - 60

2P + 2.5P = 300 + 60

4.5P = 360

360 P= 4.5

P=$ 80

Q = 300 - 2.5 * 80 = 100 units

Equilibrium quantity = 100 units

equilibrium price = $ 80 per unit

c. Refer the attached picture below for the demand and supply curve

Price Demand Supply 20 40 60 80 140 160 180 200 100 120 Qunatity

d. New Supply Curve P = 0.5Qs + 75

Qs = 2P - 150

QST 150 105 125 135 Price QD $ 50 175 $ 55 162.5 $ 60 65 137.5 70 125 75 112.5 80 100 87.5 75 95 62.5 $ 100 50 $ 105 37.5 $ 1

Equilibrium we can determine by equating the demand curve to new supply curve

300 – 2.5P = 2P - 15

2P + 2.5P = 300 + 15

4.5P = 315

P- 315 4.5 = 70

Q' = 300 - 2.5 * 70 = 125

New equilibrium quantity = 125 unit

new equilibrium price = $ 70 per unit

Refer the attched picture for the graph

$110 Price - Demand Supply --New Supply 25 50 75 175 200 225 100 125 150 Qunatity

e. The supply has increased this is due to shift in the supply curve. The supply curve has shifted to its right that is the supply has increased it may happen when

  • The cost of input decreases
  • due to technological advancement the efficiency of the existing firms improve
  • when more number of firms enter into the business then also the supply curve shifts to its right.

f. When a price floor at $ 80 is introduced then

$120 $110 Surplus Price - Demand --New Supply - Price Floor 25 50 75 175 200 225 100 125 150 Qunatity

As we can see from the above picture the quanity demanded at the price floor is 100 units but the quantity supplied at this price is equal to 145 units.

A price floor is effective when it is set above the equilibrium price, as is the case here.

Due to price floor a surplus occurs that is the scenario where quantity supplied is greater than quantity demanded. Surplus = Supply - Demand

Surplus = 145 - 100 = 45 units.

Please contact if having any query will be obliged to you for your generous support. Please help me it mean a lot to me, please help. Thank you.